Data obtained from a test of Nicorette gum a chewing gum to help smokers
If two different subjects are randomly sedected, without replacement find the probability that they are both from the placebo group.
mouth or throat soreness Nicorette (41) placebo (35)
no mouth or throat soreness Nicorette (109) placebo (118)

Answer by stanbon(75887)

Data obtained from a test of Nicorette gum a chewing gum to help smokers
If two different subjects are randomly sedected, without replacement find the probability that they are both from the placebo group.
mouth or throat soreness Nicorette (41) placebo (35)
no mouth or throat soreness Nicorette (109) placebo (118)
-----------
# in Placebo group = 35+118=153
Total number= 303
Prob(two from Placebo group)= 153C2/303C2=0.254
---------
Another way to do it is
(153/303)(152/302)=0.254
---------
Cheers,
Stan H.